# ZeroFlow: Overcoming Catastrophic Forgetting is Easier than You Think

### v2

@@ 4.1 Benchmark Setups
Forgetting scenarios, schemes, and models. We assigned the evaluation to a common catastrophic forgetting scenario, which is continual learning. We focus on four fine-tuning tasks, considering their complexity from the typical dataset CIFAR and CUB to dataset ImageNet-A, and OmniBenchmark, which has a large domain gap. For the method used on these tasks, we explore two SOTA schemes: EASE and APER. Both models are initially configured with ViT-B/16-IN21K and subsequently fine-tuned on four datasets. Following [66], each dataset is split equally into 10 tasks by class. For example, OmniBenchmark contains 300 classes, with 30 classes learned in each incremental stage. No memory is allowed for storing past examples.
Benchmark setup and details. To evaluate the application of ZeroFlow in forgetting scenarios, we include the methods described in Sec. 3.1, specifically ZO [19, 38], Sign [34], and Conserve [27, 64], in comparison with their FO counterparts using SGD and Adam optimizers [7]. Additionally, as highlighted in [64], Forward-Grad [44] which relies on forward mode automatic differentiation, potentially becomes a missing but competitive forward pass baseline. In a nutshell, ZeroFlow covers 7 forward pass methods: ZO-SGD, ZO-SGD-Sign, ZO-SGD-Conserve, ZO-Adam, ZO-Adam-Sign, ZO-Adam-Conserve, Forward-Grad.
@@
Table 2: The Memory Cost (in GB) and the Runtime Cost (in seconds) of Each Optimizer on Four Forgetting Scenarios. The per-epoch runtime in seconds (s). ZO-SGD w/ the query budget $q=1,4$ and all other optimizers w/ the query budget $q=1$.
Optimizer Memory $\Downarrow$ CIFAR-100 CUB ImageNet-A OmniBenchmark
FO-SGD 12.08 GB 59.3s 16.1s 12.2s 109.6s
ZO-SGD ($q=1$) 2.41 GB 32.4s 8.3s 6.8s 66.8s
ZO-SGD ($q=4$) 2.41 GB 111.7s 28.7s 18.0s 221.3s
ZO-SGD-Sign 2.41 GB 32.4s 8.3s 6.8s 66.8s
ZO-SGD-Conserve 2.41 GB 70.1s 15.7s 12.4s 133.7s
Forward-Grad 3.94 GB 45.9s 11.1s 9.0s 82.0s
@@
Figure 4: ZeroFlow Evaluation Results for Forgetting. We visualize the evaluation results of 2 models in the last-task accuracy.
4.2 Evaluation Results of ZeroFlow
ZeroFlow evaluation on continual learning. In Table 1, we evaluate the performance of different BP-free and BP-based (FO-SGD and FO-Adam) methods in a typical forgetting scenario (continual learning). We use two SOTA models as examples and investigate SGD and Adam optimizers, seven forward pass methods, and four commonly used datasets. Several observations are listed below,
First, the performance of the ZO method is comparable to or even surpasses that of the FO method across almost all forgetting metrics and datasets. However, as will be shown later, the FO method requires significantly more memory overhead. This suggests that forward passes alone can effectively mitigate forgetting, and the ZO method offers a simpler, more efficient alternative. In some cases, such as with ZO-Adam and ZO-SGD on OmniBenchmark, ZO methods even outperform FO methods.
Second, the Forward Grad method demonstrates competitive performance when compared to other ZO and FO methods. Unlike typical ZO methods, Forward Grad utilizes a unique forward pass mechanism, making it a promising baseline for future studies. A more intuitive trend in overcoming forgetting can be observed in Figure 4. These observations motivate further exploration into the effectiveness of the ZO method.
@@ 6 New Enhancement to Mitigate Forgetting
In this section, we further investigate several techniques to stabilize the updates for mitigating forgetting. In ZO optimization, the estimation of the gradients relies on a finite difference of the objective function (i.e., the gradients are estimated by querying the function values). We set $q=1$ in $q$-SPSA for efficiency in the previous benchmark settings. However, limited queries cannot capture the accurate ZO directions, leading to large variances in the gradient estimation. In continual learning scenarios where models learn tasks sequentially, the high variance inherent in ZO gradient estimation poses a critical challenge. This uncertainty, coupled with large parameter updates, can significantly amplify catastrophic forgetting through undesirable perturbations to previously learned knowledge. This issue is amplified for larger models such as ViTs and model-expansion architectures, due to their high-dimensional parameter spaces. Though increasing query numbers can stabilize the gradient estimates, it leads to prohibitive computational costs in continual learning settings. Thus, exploring variance-reduced optimization algorithms is crucial for practical ZO-based continual learning. Specifically, we propose two complementary approaches to stabilize the ZO optimization process while maintaining computational efficiency:
@@
Figure 7: The Variation in Function Values of the Forward Passes. Function values for new tasks is highlighted in red, old tasks is highlighted in green.
First, leveraging historical gradient information to overcome forgetting. When learning new tasks, models leverage previously learned parameters while prioritizing the preservation of crucial parameters for old tasks. To mitigate interference from new tasks, we propose reweighting old task gradients with historical gradients, which can stabilize perturbations caused by low query loops in ZO optimization. Figure 7 illustrates the function value trajectories for both old and new tasks. While FO optimization shows smooth convergence toward the global optimum, ZO optimization exhibits a more volatile path. Notably, objectives related to old tasks demonstrate smaller changes in both magnitude and variance. This observation motivates us to stabilize the optimization by reducing changes to old gradients through a linear combination with historical gradients: $g_{old}=(1-\alpha)g_{old}+\alpha g_{historical}$, where larger $\alpha$ indicates greater reliance on historical information for stability, at the cost of reduced contrast with new task gradients.
Table 3 validates the effectiveness of historical estimation in mitigating catastrophic forgetting. Modest proportions of historical information (e.g., 20%, 40%, 60%)) outperform ZO-SGD (0%), effectively controlling perturbations while maintaining a low query budget ($q=1$). While this approach helps stabilize updates, the inherent noise in ZO estimation requires additional consideration, leading us to explore complementary variance reduction techniques.
Table 3: Effectiveness of Historical Estimation in Mitigating Forgetting. The proportion of 0% denotes that the plain optimizer ZO-SGD.
@@
Table 4: Effectiveness of Sparsity-induced Estimation in Overcoming Forgetting. The proportion of 0% denotes that the plain optimizer ZO-SGD.
Metrics 0% 10% 20% 30% 40% 50% 60% 70% 80% 90%
@@
Second, sparsity-induced estimation helps to overcome forgetting. Gradient sparsification reduces the computational cost and memory usage by setting parts of the gradients to zero or by retaining a subset of the gradient components. In ZO optimization, the gradients for new tasks are often highly uncertain due to the approximation nature of the gradient estimation. To reduce this variance, we implement random sparsification by creating a seed-based mask and setting gradients outside the mask to zero. By reducing the number of non-zero gradient components, we aim to stabilize the optimization process and mitigate the noise in gradient updates. However, it is important to balance sparsity carefully, as excessively high sparsity could impair the learning of new tasks and exacerbate catastrophic forgetting, especially when too many important gradient components are zeroed out. The sparsity level is varied in our experiments, ranging from 10% to 90%.
Table 4 shows the performance of sparsity-based ZO optimization in overcoming catastrophic forgetting. We observe that the sparse technique improves the average and last accuracy across all scales, which implies that forgetting is effectively controlled. The reduction in volatility can be attributed to the sparse strategy yielding smoother gradient estimates compared to plain ZO-SGD, effectively bounding variance to a low level and thus mitigating forgetting. Moreover, the robust performance across different sparsity ratios provides strong evidence for the efficacy of variance control in addressing forgetting.

### v3

@@ 4.1 Benchmark Setups
Forgetting scenarios, schemes, and models. We assigned the evaluation to a common catastrophic forgetting scenario, which is continual learning. We focus on four fine-tuning tasks, considering their complexity from the typical dataset CIFAR and CUB to dataset ImageNet-A, and OmniBenchmark, which has a large domain gap. For the method used on these tasks, we explore two SOTA schemes: EASE and APER. Both models are initially configured with ViT-B/16-IN21K and subsequently fine-tuned on four datasets. Following (Zhou et al., 2023a), each dataset is split equally into 10 tasks by class. For example, OmniBenchmark contains 300 classes, with 30 classes learned in each incremental stage. No memory is allowed for storing past examples.
Benchmark setup and details. To evaluate the application of ZeroFlow in forgetting scenarios, we include the methods described in Section 3.1, specifically ZO (Ghadimi & Lan, 2013; Malladi et al., 2023), Sign (Liu et al., 2019), and Conserve (Kim et al., 2021; Zhang et al., 2024), in comparison with their FO counterparts using SGD and Adam optimizers (Chen et al., 2019). Additionally, as highlighted in (Zhang et al., 2024), Forward-Grad (Ren et al., 2022) which relies on forward mode automatic differentiation, potentially becomes a missing but competitive forward pass baseline. In a nutshell, ZeroFlow covers 7 forward pass methods: ZO-SGD, ZO-SGD-Sign, ZO-SGD-Conserve, ZO-Adam, ZO-Adam-Sign, ZO-Adam-Conserve, Forward-Grad.
@@
Table 2: Memory Cost (GB) and Runtime Cost (s) of Each Optimizer on 3 Forgetting Scenarios. The per-epoch runtime in seconds (s). ZO-SGD w/ query budget $q=1,4$ and all other optimizers w/ query budget $q=1$.
Optimizer Memory $\Downarrow$ CIFAR-100 CUB ImageNet-A
FO-SGD 12.08 GB 59.3s 16.1s 12.2s
ZO-SGD ($q=1$) 2.41 GB 32.4s 8.3s 6.8s
ZO-SGD ($q=4$) 2.41 GB 111.7s 28.7s 18.0s
ZO-SGD-Sign 2.41 GB 32.4s 8.3s 6.8s
ZO-SGD-Conserve 2.41 GB 70.1s 15.7s 12.4s
Forward-Grad 3.94 GB 45.9s 11.1s 9.0s
4.2 Evaluation Results of ZeroFlow
ZeroFlow evaluation on continual learning. In Table 1, we evaluate the performance of different BP-free and BP-based (FO-SGD and FO-Adam) methods in a typical forgetting scenario (continual learning). We use two SOTA models as examples (EASE (Zhou et al., 2024b) and APER (Zhou et al., 2023a)) and investigate SGD and Adam optimizers, seven forward pass methods, and four commonly used datasets. Several observations are listed below,
First, the performance of ZO method is comparable to or even surpasses that of the FO method across almost all forgetting metrics and datasets. However, as will be shown later, the FO method requires significantly more memory overhead. This suggests that forward passes alone can effectively mitigate forgetting, and the ZO method offers a simpler, more efficient alternative. In some cases, such as with ZO-Adam and ZO-SGD on OmniBenchmark, ZO methods even outperform FO methods.
Second, Forward Grad demonstrates competitive performance when compared to other ZO and FO methods. Unlike typical ZO methods, Forward Grad utilizes a unique forward pass mechanism, making it a promising baseline for future studies. A more intuitive trend in overcoming forgetting refer to Figure 4. These observations motivate further exploration into the effectiveness of ZO method.
@@ 4.2 Evaluation Results of ZeroFlow
Figure 4: ZeroFlow Evaluation Results for Forgetting. We visualize the evaluation of 2 models in last-task accuracy.
@@ 6 New Enhancement to Mitigate Forgetting
In ZO optimization, the estimation of the gradients relies on a finite difference of the objective function. We set query budget $q=1$ in the benchmark for efficiency. However, limited queries cannot capture the accurate ZO directions. When the model learns tasks sequentially, the high variance inherent in ZO gradient estimation poses a critical challenge. Though increasing query numbers can stabilize the gradient estimates, it leads to prohibitive overhead Thus, exploring variance-reduced optimization algorithms is crucial for ZO-based CL. Specifically, we propose 3 enhancements to stabilize the ZO optimization process:
Figure 7: Effectiveness of Hybrid ZO in Overcoming Forgetting. In Hybrid ZO, backward benefits from forward pass.
Enhancement 1: Hybrid ZO to overcome forgetting. While ZO methods does not explicitly minimize sharpness, it stabilizes optimization by approximating gradients and assessing the rate of change in loss function through perturbations. This indirect approach helps reduce the curvature of the loss landscape, steering the optimization away from sharp and unstable regions. This insight motivates us to investigate Hybrid ZO method. Figure 7 illustrates results hybrid ZO. We first use FO to coarsely optimize to a local minimum (first 140 or 160 epochs) and then refine the solution by searching for flatter regions around it using ZO (last 30 or 60 epochs). As the first two subfigures in Figure 7, ZO provides only limited gains to FO. This is because FO inherits strong generalization from the pretrained backbone but loses its generalization ability quickly after two incremental stages. In later stages, ZO helps to remedy the vulnerabilities of backbone trained by FO, leading to significant enhancements compared to the FO baseline.
@@
Figure 8: Variation in Function Values of Forward Passes. Function values for new tasks is highlighted in red, old tasks is highlighted in green.
Enhancement 2: Leverage historical information to overcome forgetting. When learning new tasks, models leverage previously learned parameters while prioritizing the preservation of crucial parameters for old tasks. To mitigate interference from new tasks, we propose reweighting old task gradients with historical gradients, which can stabilize perturbations caused by low query loops in ZO optimization. Figure 8 illustrates the function value trajectories for both old and new tasks. While FO optimization shows smooth convergence toward the global optimum, ZO optimization exhibits a more volatile path. Notably, objectives related to old tasks demonstrate smaller changes in both magnitude and variance. This observation motivates us to stabilize the optimization by reducing changes to old gradients through a linear combination with historical gradients: $g_{old}=(1-\alpha)g_{old}+\alpha g_{historical}$, where larger $\alpha$ indicates greater reliance on historical information for stability, at the cost of reduced contrast with new task gradients.
Table 3: Effectiveness of Historical Estimation in Mitigating Forgetting. Proportion of 0% denotes that the plain optimizer ZO-SGD.
@@
In Table 3, we validate the effectiveness of historical estimation in mitigating catastrophic forgetting. Modest proportions of historical information (e.g., 20%, 40%, 60%)) outperform ZO-SGD (0%), effectively controlling perturbations while maintaining a low query budget ($q=1$).
Enhancement 3: Sparsity-induced estimation helps to overcome forgetting. In ZO optimization, the gradients for new tasks are often highly uncertain due to the approximation nature of the gradient estimation. To reduce this variance, we implement random sparsification by creating a seed-based mask and setting gradients outside the mask to zero. By reducing the number of non-zero gradient components, we aim to stabilize the optimization process and mitigate the noise in gradient updates.
Table 4: Effectiveness of Sparsity-induced Estimation in Overcoming Forgetting. Proportion of 0% denotes the plain ZO-SGD.
Ratio 0% 10% 20% 30% 40% 50% 60% 70% 80% 90%
@@
In Table 4, we report the performance of sparsity-induced ZO in overcoming forgetting. The sparsity level is varied in this experiments, ranging from 10% to 90%. We observe that the sparse technique improves the average and last accuracy across all scales, which implies that forgetting is effectively controlled. The reduction in volatility can be attributed to the sparse strategy yielding smoother gradient estimates compared to plain ZO-SGD, effectively bounding variance to a low level and thus mitigating forgetting. Moreover, the robust performance across different sparsity ratios provides strong evidence for the efficacy of variance control in addressing forgetting.
@@ 7 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.